# Tight Bounds for Schrödinger Potential Estimation in Unpaired Data Translation

### v2

@@ Appendix D Log-potential variance and conditional variance bounds
$M\leqslant-\mathbb{E}\varphi^{*}(X_{T}^{*})\leqslant L\mathbb{E}\left\|\Sigma^{-1/2}(X_{T}^{*}-m)\right\|^{2}+M=L\mathrm{Tr}\big(\Sigma^{-1}\mathrm{Var}(X_{T}^{*})\big)+L\left\|\Sigma^{-1/2}(\mathbb{E}X_{T}^{*}-m)\right\|^{2}+M.$
@@
$\mathrm{Var}\left(\mathbb{E}\left[\widetilde{\varphi}(X_{T}^{*})-\widetilde{\varphi}^{*}(X_{T}^{*})\,\big|\,X_{0}\right]\right)$ $\leqslant 2\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}\big[(\widetilde{\varphi}-\widetilde{\varphi}^{*})_{+}e^{\varphi^{*}}\big](X_{0})}{{\mathcal{T}}_{T}\big[e^{\varphi^{*}}\big](X_{0})}\right)+2\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}\big[(\widetilde{\varphi}-\widetilde{\varphi}^{*})_{-}e^{\varphi^{*}}\big](X_{0})}{{\mathcal{T}}_{T}\big[e^{\varphi^{*}}\big](X_{0})}\right).$
An upper bound on the terms in the right-hand side easily follows from Lemma D.3. Indeed, introducing
@@
and applying Lemma D.3, we obtain that
$\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}[(\widetilde{\varphi}-\widetilde{\varphi}^{*})_{+}e^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)$ $\leqslant\varphi_{\max}^{2}\left(\frac{{\mathcal{T}}_{\infty}[(\widetilde{\varphi}-\widetilde{\varphi}^{*})_{+}e^{\varphi^{*}}]}{\varphi_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{2/{\mathcal{K}}(T)}\exp\left\{2M\big({\mathcal{K}}(T)-1\big)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
$\quad+\frac{\varphi_{\max}^{2}}{e^{2}}\left({\mathcal{K}}(T)-\frac{1}{{\mathcal{K}}(T)}\right)^{2}\left(\frac{{\mathcal{T}}_{\infty}[(\widetilde{\varphi}-\widetilde{\varphi}^{*})_{+}e^{\varphi^{*}}]}{\varphi_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{4/{\mathcal{K}}(T)-2{\mathcal{K}}(T)}$
@@
$\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}[(\widetilde{\varphi}-\widetilde{\varphi}^{*})_{-}e^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)$ $\leqslant\varphi_{\max}^{2}\left(\frac{{\mathcal{T}}_{\infty}[(\widetilde{\varphi}-\widetilde{\varphi}^{*})_{-}e^{\varphi^{*}}]}{\varphi_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{2/{\mathcal{K}}(T)}\exp\left\{2M\big({\mathcal{K}}(T)-1\big)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
@@
Applying Lemma F.10 and using the bound
@@
$\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)$ $=\mathbb{E}\,\frac{{\mathcal{T}}_{T}[(\widetilde{\varphi}-\widetilde{\varphi}^{*})^{2}e^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}$
@@
$\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)$ $\geqslant H_{\max}^{1-{\mathcal{K}}(T)}\left(\frac{{\mathcal{T}}_{\infty}[(\widetilde{\varphi}-\widetilde{\varphi}^{*})^{2}e^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{{\mathcal{K}}(T)}\exp\left\{-M\left(1-\frac{1}{{\mathcal{K}}(T)}\right)-\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\},$
@@
$\leqslant H_{\max}^{{\mathcal{K}}(T)-1}\,\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)\,\exp\left\{M\left(1-\frac{1}{{\mathcal{K}}(T)}\right)+\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}.$
@@
$\quad+\frac{2\varphi_{\max}^{2+2{\mathcal{K}}(T)-4/{\mathcal{K}}(T)}H_{\max}^{({\mathcal{K}}(T)-1)(2/{\mathcal{K}}^{2}(T)-1)}}{e^{2}}\left({\mathcal{K}}(T)-\frac{1}{{\mathcal{K}}(T)}\right)^{2}\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)^{2/{\mathcal{K}}^{2}(T)-1}$
@@
Then, due to Lemma F.10, for any $x\in\mathbb{R}^{d}$, it holds that
$\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](x)}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](x)}\leqslant\big(F(x)\big)^{1-1/{\mathcal{K}}(T)}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{1/{\mathcal{K}}(T)}\exp\left\{M\big({\mathcal{K}}(T)-1\big)+{\mathcal{A}}(x,T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
@@
$\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](x)}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](x)}\geqslant\big(F(x)\big)^{1-{\mathcal{K}}(T)}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{{\mathcal{K}}(T)}\exp\left\{-M\left(1-\frac{1}{{\mathcal{K}}(T)}\right)-{\mathcal{A}}(x,T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\},$
@@
$\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\leqslant\big(F_{\max}\big)^{1-1/{\mathcal{K}}(T)}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{1/{\mathcal{K}}(T)}\exp\left\{M\big({\mathcal{K}}(T)-1\big)+\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
@@
$\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\geqslant\big(F_{\max}\big)^{1-{\mathcal{K}}(T)}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{{\mathcal{K}}(T)}\exp\left\{-M\left(1-\frac{1}{{\mathcal{K}}(T)}\right)-\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
@@
$\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)$ $\leqslant\frac{1}{4}\Bigg(\big(F_{\max}\big)^{1-1/{\mathcal{K}}(T)}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{1/{\mathcal{K}}(T)}\exp\left\{M\big({\mathcal{K}}(T)-1\big)+\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
$\quad-\big(F_{\max}\big)^{1-{\mathcal{K}}(T)}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{{\mathcal{K}}(T)}\exp\left\{-M\left(1-\frac{1}{{\mathcal{K}}(T)}\right)-\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}\Bigg)^{2}.$
@@
$\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)$ $\leqslant\frac{F_{\max}^{2}}{4}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{F_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{2/{\mathcal{K}}(T)}\exp\left\{2M\big({\mathcal{K}}(T)-1\big)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
$\quad\cdot\left(1-\exp\left\{-\left({\mathcal{K}}(T)-\frac{1}{{\mathcal{K}}(T)}\right)\left(M+\log\frac{F_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}\right)-2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}\right)^{2}.$
@@
$\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)$ $\leqslant\frac{F_{\max}^{2}}{4}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{F_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{2/{\mathcal{K}}(T)}\exp\left\{2M\big({\mathcal{K}}(T)-1\big)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
$\quad\cdot\left[\left({\mathcal{K}}(T)-\frac{1}{{\mathcal{K}}(T)}\right)\left(M+\log\frac{F_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}\right)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right]^{2}.$
@@
$\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)$ $\leqslant F_{\max}^{2}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{F_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{2/{\mathcal{K}}(T)}\exp\left\{2M\big({\mathcal{K}}(T)-1\big)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
$\quad\cdot\left\{\left({\mathcal{K}}(T)-\frac{1}{{\mathcal{K}}(T)}\right)^{2}\left[M^{2}+\log^{2}\left(\frac{F_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}\right)\right]+2\big(\overline{{\mathcal{A}}}(T)\big)^{2}\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)^{2}\right\}.$
@@
$\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)$ $\leqslant F_{\max}^{2}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{F_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{2/{\mathcal{K}}(T)}\exp\left\{2M\big({\mathcal{K}}(T)-1\big)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
@@
We would like to note that, according to Lemma F.8, ${\mathsf{KL}}(\pi^{*},\pi^{\varphi})=\mathbb{E}\zeta\leqslant 2\|\zeta\|_{\psi_{1}}$. This implies that $\varepsilon\in(0,1)$. Then, applying Hölder’s inequality, we obtain that
@@
Let us elaborate on $\mathbb{E}\zeta^{(1+\varepsilon)/\varepsilon}$. Using Lemma F.8 again, we observe that the expectation of $\zeta^{(1+\varepsilon)/\varepsilon}$ does not exceed
@@
$\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)\leqslant 2\mathrm{Var}\left(\mathbb{E}\left[\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\,\big|\,X_{0}\right]\right)+4e^{3/2}\,{\mathsf{KL}}(\pi^{*},\pi^{\varphi})\|\zeta\|_{\psi_{1}}\left(1+\frac{1}{2}\log\frac{2e\|\zeta\|_{\psi_{1}}}{{\mathsf{KL}}(\pi^{*},\pi^{\varphi})}\right).$

### v3

@@ Appendix C Log-potential variance and conditional variance bounds
$-M\leqslant-\mathbb{E}\varphi^{*}(X_{T}^{*})$ $\leqslant L\mathbb{E}\left\|\Sigma^{-1/2}(X_{T}^{*}-m)\right\|^{2}+M$
$=L\mathrm{Tr}\big(\Sigma^{-1}\mathrm{Var}(X_{T}^{*})\big)+L\left\|\Sigma^{-1/2}(\mathbb{E}X_{T}^{*}-m)\right\|^{2}+M.$
@@
$\mathrm{Var}\left(\mathbb{E}\left[\widetilde{\varphi}(X_{T}^{*})-\widetilde{\varphi}^{*}(X_{T}^{*})\,\big|\,X_{0}\right]\right)$ $\leqslant 2\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}\big[(\widetilde{\varphi}-\widetilde{\varphi}^{*})_{+}e^{\varphi^{*}}\big](X_{0})}{{\mathcal{T}}_{T}\big[e^{\varphi^{*}}\big](X_{0})}\right)$
$\quad+2\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}\big[(\widetilde{\varphi}-\widetilde{\varphi}^{*})_{-}e^{\varphi^{*}}\big](X_{0})}{{\mathcal{T}}_{T}\big[e^{\varphi^{*}}\big](X_{0})}\right).$
An upper bound on the terms in the right-hand side easily follows from Lemma C.3. Indeed, introducing
@@
and applying Lemma C.3, we obtain that
$\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}[(\widetilde{\varphi}-\widetilde{\varphi}^{*})_{+}e^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)$ $\leqslant\varphi_{\max}^{2}\left(\frac{{\mathcal{T}}_{\infty}[(\widetilde{\varphi}-\widetilde{\varphi}^{*})_{+}e^{\varphi^{*}}]}{\varphi_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{2/{\mathcal{K}}(T)}$
$\cdot\exp\left\{2M\big({\mathcal{K}}(T)-1\big)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
$\cdot\left\{M^{2}\left({\mathcal{K}}(T)-\frac{1}{{\mathcal{K}}(T)}\right)^{2}+2\big(\overline{{\mathcal{A}}}(T)\big)^{2}\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)^{2}\right\}$
$+\left({\mathcal{K}}(T)-\frac{1}{{\mathcal{K}}(T)}\right)^{2}\left(\frac{{\mathcal{T}}_{\infty}[(\widetilde{\varphi}-\widetilde{\varphi}^{*})_{+}e^{\varphi^{*}}]}{\varphi_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{4/{\mathcal{K}}(T)-2{\mathcal{K}}(T)}$
$\cdot\frac{\varphi_{\max}^{2}}{e^{2}}\cdot\exp\left\{2M\big({\mathcal{K}}(T)-1\big)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\},$
@@
$\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}[(\widetilde{\varphi}-\widetilde{\varphi}^{*})_{-}e^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)$
$\leqslant\varphi_{\max}^{2}\left(\frac{{\mathcal{T}}_{\infty}[(\widetilde{\varphi}-\widetilde{\varphi}^{*})_{-}e^{\varphi^{*}}]}{\varphi_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{2/{\mathcal{K}}(T)}\exp\left\{2M\big({\mathcal{K}}(T)-1\big)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
@@
Applying Lemma E.10 and using the bound
@@
$\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)$
$=\mathbb{E}\,\frac{{\mathcal{T}}_{T}[(\widetilde{\varphi}-\widetilde{\varphi}^{*})^{2}e^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}$
@@
$\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)$ $\geqslant H_{\max}^{1-{\mathcal{K}}(T)}\left(\frac{{\mathcal{T}}_{\infty}[(\widetilde{\varphi}-\widetilde{\varphi}^{*})^{2}e^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{{\mathcal{K}}(T)}$
$\quad\cdot\exp\left\{-M\left(1-\frac{1}{{\mathcal{K}}(T)}\right)-\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\},$
@@
$\leqslant H_{\max}^{{\mathcal{K}}(T)-1}\,\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)$
$\quad\cdot\exp\left\{M\left(1-\frac{1}{{\mathcal{K}}(T)}\right)+\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}.$
@@
$\quad+\frac{2\varphi_{\max}^{2+2{\mathcal{K}}(T)-4/{\mathcal{K}}(T)}H_{\max}^{({\mathcal{K}}(T)-1)(2/{\mathcal{K}}^{2}(T)-1)}}{e^{2}}$
$\quad\cdot\left({\mathcal{K}}(T)-\frac{1}{{\mathcal{K}}(T)}\right)^{2}\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)^{2/{\mathcal{K}}^{2}(T)-1}$
@@
Then, due to Lemma E.10, for any $x\in\mathbb{R}^{d}$, it holds that
$\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](x)}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](x)}$ $\leqslant\big(F(x)\big)^{1-1/{\mathcal{K}}(T)}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{1/{\mathcal{K}}(T)}$
$\quad\cdot\exp\left\{M\big({\mathcal{K}}(T)-1\big)+{\mathcal{A}}(x,T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
@@
$\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](x)}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](x)}$ $\geqslant\big(F(x)\big)^{1-{\mathcal{K}}(T)}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{{\mathcal{K}}(T)}$
$\quad\cdot\exp\left\{-M\left(1-\frac{1}{{\mathcal{K}}(T)}\right)-{\mathcal{A}}(x,T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\},$
@@
$\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}$ $\leqslant\big(F_{\max}\big)^{1-1/{\mathcal{K}}(T)}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{1/{\mathcal{K}}(T)}$
$\quad\cdot\exp\left\{M\big({\mathcal{K}}(T)-1\big)+\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
@@
$\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}$ $\geqslant\big(F_{\max}\big)^{1-{\mathcal{K}}(T)}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{{\mathcal{K}}(T)}$
$\quad\cdot\exp\left\{-M\left(1-\frac{1}{{\mathcal{K}}(T)}\right)-\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
@@
$\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)$ $\leqslant\frac{1}{4}\Bigg(\big(F_{\max}\big)^{1-1/{\mathcal{K}}(T)}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{1/{\mathcal{K}}(T)}$
$\quad\cdot\exp\left\{M\big({\mathcal{K}}(T)-1\big)+\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
$\quad-\big(F_{\max}\big)^{1-{\mathcal{K}}(T)}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{{\mathcal{K}}(T)}$
$\quad\cdot\exp\left\{-M\left(1-\frac{1}{{\mathcal{K}}(T)}\right)-\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}\Bigg)^{2}.$
@@
$\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)$
$\leqslant\frac{F_{\max}^{2}}{4}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{F_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{2/{\mathcal{K}}(T)}\exp\left\{2M\big({\mathcal{K}}(T)-1\big)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
$\cdot\left(1-\exp\left\{-\left({\mathcal{K}}(T)-\frac{1}{{\mathcal{K}}(T)}\right)\left(M+\log\frac{F_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}\right)-2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}\right)^{2}.$
@@
$\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)$
$\leqslant\frac{F_{\max}^{2}}{4}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{F_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{2/{\mathcal{K}}(T)}\exp\left\{2M\big({\mathcal{K}}(T)-1\big)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
$\cdot\left[\left({\mathcal{K}}(T)-\frac{1}{{\mathcal{K}}(T)}\right)\left(M+\log\frac{F_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}\right)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right]^{2}.$
@@
$\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)$
$\leqslant F_{\max}^{2}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{F_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{2/{\mathcal{K}}(T)}\exp\left\{2M\big({\mathcal{K}}(T)-1\big)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
$\cdot\left\{\left({\mathcal{K}}(T)-\frac{1}{{\mathcal{K}}(T)}\right)^{2}\left[M^{2}+\log^{2}\left(\frac{F_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}\right)\right]+2\big(\overline{{\mathcal{A}}}(T)\big)^{2}\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)^{2}\right\}.$
@@
$\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)$ $\leqslant F_{\max}^{2}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{F_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{2/{\mathcal{K}}(T)}$
$\quad\cdot\exp\left\{2M\big({\mathcal{K}}(T)-1\big)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
@@
We would like to note that, according to Lemma E.8, ${\mathsf{KL}}(\pi^{*},\pi^{\varphi})=\mathbb{E}\zeta\leqslant 2\|\zeta\|_{\psi_{1}}$. This implies that $\varepsilon\in(0,1)$. Then, applying Hölder’s inequality, we obtain that
@@
Let us elaborate on $\mathbb{E}\zeta^{(1+\varepsilon)/\varepsilon}$. Using Lemma E.8 again, we observe that the expectation of $\zeta^{(1+\varepsilon)/\varepsilon}$ does not exceed